# Dirichlet process mixtures of block $g$ priors for model selection and prediction in linear models

### v1

@@ 4.6 Intrinsic consistency
$f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid a,b,\boldsymbol{\gamma})=\int\phi\left(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\boldsymbol{0},\sigma^{2}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{\Lambda}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right)\\ f\left(\tilde{\boldsymbol{g}}\mid\rho,\gamma,\tau^{2}=n,a,b\right)f(\rho\mid\boldsymbol{\gamma},\alpha)f(\alpha\mid\boldsymbol{\gamma})\mathop{}\!\mathrm{d}\tilde{g}_{1}\ldots\mathop{}\!\mathrm{d}\tilde{g}_{p_{\boldsymbol{\gamma}}}\mathop{}\!\mathrm{d}\rho\mathop{}\!\mathrm{d}\alpha,$
The proof is straightforward and relies on the ability to swap the integration and limit operations.
@@ 5 Computation
In the case of linear models (as well as some classes of generalized linear models that admit similar representations after introducing latent variables, such as logistic regression models, Polson et al., 2013), it is possible to construct MCMC algorithms for model selection under the DP mixtures of block $g$ prior that require very minimal tuning. To do so, we take advantage of the conditional conjugacy of the priors and, when possible, we integrate out the intercept $\beta_{0}$, the vector of regression coefficients $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ and/or the variance $\sigma^{2}$ when deriving conditional posteriors. Additionally, we represent the shrinkage coefficients $g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}}$ in terms of their unique values $\tilde{\boldsymbol{g}}_{\boldsymbol{\gamma}}=(\tilde{g}_{1},\ldots,\tilde{g}_{K_{\boldsymbol{\gamma}}})$ and the group indicators $\boldsymbol{\xi}_{\boldsymbol{\gamma}}=(\xi_{1},\ldots,\xi_{p_{\boldsymbol{\gamma}}})$ (recall Section 3). The resulting posterior takes the form
@@
where $f(\boldsymbol{y}\mid\boldsymbol{\gamma},\tilde{g}_{1},\ldots,\tilde{g}_{K_{\boldsymbol{\gamma}}},\xi_{1},\ldots,\xi_{p_{\boldsymbol{\gamma}}})$ corresponds to (3.2) with $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$, and $p(\boldsymbol{\gamma})$ is an appropriate prior on the space of models, e.g., a Beta-Binomial prior
$f(\boldsymbol{\gamma})=\frac{\Gamma(c+d)}{\Gamma(c)\Gamma(d)}\frac{\Gamma(c+p_{\boldsymbol{\gamma}})\Gamma(d+p-p_{\boldsymbol{\gamma}})}{\Gamma(c+d+p)}.$
Our MCMC algorithm then alternates sampling from the full conditionals $f(\boldsymbol{\gamma},\tilde{\boldsymbol{g}},\boldsymbol{\xi}\mid\cdots)$, $f(\boldsymbol{\xi}\mid\cdots)$, $f(\alpha\mid\cdots)$, $f(\tilde{\boldsymbol{g}}\mid\cdots)$, $f(\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\cdots)$ and $f(\sigma^{2}\mid\cdots)$. In particular, to sample from $f(\boldsymbol{\gamma},\tilde{\boldsymbol{g}},\boldsymbol{\xi}\mid\cdots)$, we use a random walk Metropolis algorithm in which, at each iteration, we propose to either add one variable, remove one variable, or swap one variable currently in the model with one that is not. If a variable is added to the model, the corresponding value of $\xi_{i}$ and, if necessary, a new value of $\tilde{g}_{k}$, are proposed from the prior distributions on these parameters. This is technically a Reversible Jump MCMC step (Green, 1995), albeit a very simple one. In spite of this simplicity, the algorithm seems to perform quite well. On the other hand, once we condition on the model, sampling from $f(\boldsymbol{\xi}\mid\cdots)$ can be accomplished using any of the collapsed samplers for non-conjugate Dirichlet process mixture models (e.g., see Neal, 2000). To sample $f(\tilde{\boldsymbol{g}}\mid\cdots)$ we resort to a slight variant of the slice sampler introduced in Liu et al. (2012). Finally, sampling from $f(\alpha\mid\cdots)$ is accomplished through the use of a random walk Metropolis-Hastings algorithm with log-Gaussian proposals. This is the only step of the algorithm that requires tuning of hyperparameters. Details of the algorithm are provided in Section F of the supplementary materials, and an implementation of the code is available from https://github.com/Anupreet-Porwal/DP-mix-block-g-prior.
@@ 6.1 Conditional Lindley paradox
Our first simulation study replicates the setting used to construct Figure 1 and shows empirical evidence supporting the theoretical results discussed in Sections 4.3 and 4.4. More concretely, we consider a total 150 simulations, each of which involves a sequence of datasets generated under model $\mathcal{M}_{a}:y_{i}=\beta_{0}+\beta_{1}x_{i,1}+\beta_{2}x_{i,2}+\epsilon_{i}$ for $i=1,\ldots,100$. All elements of a given sequence of datasets share the same values of $\beta_{0}$, $\beta_{1}$, $\boldsymbol{x}_{1}=(x_{1,1},\ldots,x_{100,1})^{\prime}$, $\boldsymbol{x}_{2}=(x_{1,2},\ldots,x_{100,2})^{\prime}$ and $\boldsymbol{\epsilon}=(\epsilon_{1},\ldots,\epsilon_{100})^{\prime}$. In particular, we set $\beta_{0}=0.5$, $\beta_{1}=1$ and generate $\epsilon$ from a standard multivariate Gaussian distribution and each of the pairs $(x_{i,1},x_{i,2})^{\prime}$ from a zero-mean bivariate Gaussian distribution with unit marginal standard deviations and correlation $\eta$. The various datasets within each sequence are then constructed by considering a grid of values for $\beta_{2}$ in the interval $[0,240]$. We are interested in the behavior of $B_{a,0}$, the Bayes factor under the DP mixture of $g$ priors that uses the hyper-$g/n$ as the baseline measure comparing the true model $\mathcal{M}_{a}$ against the simpler model $\mathcal{M}_{0}:y_{i}=\beta_{0}+\beta_{1}x_{i,1}+\epsilon_{i}$. We also investigate the behavior of $\mathsf{Pr}(\xi_{1}\neq\xi_{2}\mid\boldsymbol{y})$, the posterior probability that the model assigns different shrinkage parameters to each of the variables in the model (recall Section 4.3).
Figure 3 shows the results of this simulation study for $\eta=0$ and $\eta=0.5$. Compared to Figure 1, the curves are somewhat noisy. This is an artifact of the Monte Carlo noise introduced by our MCMC algorithm. Indeed, recall that the Bayes factor depicted in 1 is available in “closed form”. It could be mitigated by increasing the number of iterations used by our MCMC algorithm (which, for the purpose of these graphs, was set to 20,000, obtained after a burning the first 2,000 iterations and thinning the remaining ones every 15 samples). With that caveat in mind, we note that the curves for $\log\left(B_{a,0}(\boldsymbol{y})\right)$ decrease as $\beta_{2}$ increases but, unlike Figure 1, both seem to stabilize towards an asymptote. This agrees with the behavior predicted by Theorem 4.4. Similarly, and as predicted by Theorem 4.3, we can see that $\mathsf{Pr}(\xi_{1}\neq\xi_{2}\mid\boldsymbol{y})$ seems to converge to 1 as $\beta_{2}$ grows.
@@ 6.2 Model selection, estimation and prediction performance
We conducted a second simulation study to compare the model selection, estimation, and prediction performance of procedures based on DP mixtures of block $g$ priors with a hyper-$g/n$ base measure with that of competing procedures. This simulation study, which setup is inspired by Denti et al. (2023), assumes $n=500$ and considers six scenarios that arise from combining three different values for the total number of variables $p$ and two different levels of multicolinearity across the predictors. In terms of the number of variables, we consider scenarios with $p=250$, $p=500$ and $p=750$. For all three values of $p$, $100$ of the coefficients are randomly sampled from a normal distribution with mean 0 and standard deviation 10, $100$ are are randomly sampled from a standard normal distribution, and the remainder are set to 0. As in the first simulation study, the vectors of covariates associated with the each observation are generated from a zero-mean multivariate normal distribution with unit marginal variances and correlation $\eta$ across all pairs of covariates, with either $\eta=0$ or $\eta=0.5$. For each of the six scenarios, we generate 100 datasets to be used in evaluating the different procedures.
@@
In terms of competing approaches, we consider the following: (a) a standard hyper-$g/n$ mixture of $g$ priors (which corresponds to letting $\alpha\to 0$ in our method, and is labeled “g-prior” in the sequel); (b) a version of the block $g$ prior of Som (2014) with a hyper-$g/n$ hyperprior and known blocking structures where the covariates are allocated to $K=2$ groups: one made of all the covariates associated with coefficients generated from a normal distribution with standard deviation 10 plus half, randomly chosen variables associated with the zero coefficients, and another one made of the rest (labeled “Som et al. ($K=2$)”); (c) a version of the block $g$ prior of Som (2014) with $K=3$ fixed groups of covariates: one made of all the covariates associated with coefficients generated from a normal distribution with standard deviation 10, one made of the covariates associated with coefficients generated from a standard normal distribution, and one made of the covariates associated with zero coefficients (labeled “Som et al. ($K=3$)”); (d) a “global-local” $g$-prior with distinct but identically distributed shrinkage parameters for each coefficient, following a hyper-$g/n$ prior (this corresponds to $\alpha\to\infty$ in our method, and we label it “GL-g”); (e) the adaptive Lasso (ALasso) of (Huang et al., 2008). For benchmarkig estimation and prediction performance, we also compare against (f) the Horseshoe prior (Carvalho et al., 2010, labeled “Horseshoe” in the sequel) and (g) the Horseshoe-Pit prior (Denti et al., 2023, labeled “HSM”). Computation under standard mixtures of $g$ priors relies on version 1.7.1 of the R package BAS, while computation under block $g$ priors with known blocking structures relies on a slight variation of our own code for the Diriclet mixtures of $g$ priors. Computation for the adaptive Lasso relies on version 4.1.6 of the R package glmnet. Computation under the Horseshoe prior relies on version 1.2 of the R package bayesreg, while computation under the Horseshoe-Pit prior relies on code from the author of that manuscript which, at the time of this writing, is available at https://github.com/Fradenti/HorseshoeMix. For all Bayesian procedures that require a prior on model space, we assign $\boldsymbol{\gamma}$ a Beta-Binomial prior (e.g., see Scott & Berger, 2010 and Porwal & Raftery, 2022),
$f(\boldsymbol{\gamma})=\frac{\Gamma\left(1+\sum_{k=1}^{p}\gamma_{k}\right)\Gamma\left(1+p-\sum_{k=1}^{p}\gamma_{k}\right)}{\Gamma\left(2+p\right)}.$
Furthermore, in order to avoid improper priors when $p\geq n$, the prior on models is constrained so that models for which $\sum_{k=1}^{p}\gamma_{i}>p-2$ receive zero probability.
Power (“small” coeffs)
Power (“large” coeffs)
Type I error (null coeffs)
Power (“small” coeffs)
Power (“large” coeffs)
Type I error (null coeffs)
@@
$\eta=0$ $\eta=0.5$
$g$-prior 0.856 0.986 0.005 0.793 0.979 0.005
Som et al. $(K=2)$ 0.913 0.986 0.037 0.876 0.980 0.046
Som et al. $(K=3)$ 1.000 0.992 1.000 1.000 0.990 1.000
GL-$g$ 0.908 0.990 0.054 0.869 0.987 0.060
DP block-$g$ 0.904 0.989 0.045 0.867 0.988 0.056
ALasso 0.569 0.956 0.011 0.534 0.942 0.119
@@
$\eta=0$ $\eta=0.5$
$g$-prior 0.814 0.979 0.000 0.721 0.968 0.000
Som et al. $(K=2)$ 0.873 0.982 0.005 0.819 0.974 0.007
Som et al. $(K=3)$ 0.923 0.986 1.000 0.901 0.981 1.000
GL-$g$ 0.866 0.984 0.006 0.812 0.979 0.008
DP block-$g$ 0.867 0.984 0.006 0.814 0.979 0.009
ALasso 0.117 0.745 0.056 0.050 0.581 0.028
@@
$\eta=0$ $\eta=0.5$
$g$-prior 0.607 0.959 0.000 0.479 0.942 0.000
Som et al. $(K=2)$ 0.863 0.982 0.003 0.793 0.975 0.006
Som et al. $(K=3)$ 0.884 0.985 0.488 0.844 0.980 0.502
GL-$g$ 0.849 0.985 0.003 0.783 0.980 0.004
DP block-$g$ 0.852 0.985 0.003 0.789 0.980 0.006
ALasso 0.106 0.722 0.054 0.041 0.553 0.023
@@
Table 1: Power for “small” (generated from a $\mathcal{N}(0,1)$ distribution) and “large” (generated from a $\mathcal{N}(0,10)$ distribution) coefficients, and mean type I error for null coefficients ($\beta=0$) in our second simulation study. For the purpose of this table, coefficients are considered “significant” is their posterior inclusion probability is greater than 0.5.
@@
In the introduction, we motivated DP mixtures of block $g$ priors by arguing that it should lead to higher power for detecting smaller coefficients. We also claimed that procedures that pre-select a blocking of the coefficients can be very sensitive to this choice. Evidence of these claims is presented in Table 1, which shows the estimated power associated with identifying “large” (those generated from a zero-mean normal with standard deviation 10) and “small” (those generated from a standard normal distribution) coefficients, as well as the type I error (for the null coefficients) under the various procedures. For the purpose of this table, coefficients are considered “significant” if their posterior inclusion probability is greater than 0.5 (in the case of Bayesian procedures), or if they are included in the optimal model after applying generalized cross-validation to identify the optimal shrinkage parameter (for ALasso). First, we note that ALasso shows by far the lowest power to detect small signals. Even for large coefficients, the performance of ALasso degrades substantially in “large $p$” scenarios. The same was true of other penalized likelihood procedures we tried (results not shown). Next, we note that the standard $g$ prior tends to have lower power than the other Bayesian procedures we consider. As we expected, the difference is particularly pronounced for the “small” coefficients. The procedure of Som (2014) with $K=3$ tends to be the one with the highest power, but it also comes with extremely high type I error rates. This is because, by parceling out the null coefficients into a separate cluster, the block $g$ prior ends up overfitting by learning a very small shrinkage factor for the coefficients in this block. The results for Som (2014) with $K=2$ indicate that this issue can be addressed by having the null coefficients assigned to blocks that contain some significant coefficients. This solution is, however, clearly impractical in real applications, as we do not know which coefficients are likely to be significant in the first place. Finally, we note that the performance of GL-$g$ and DP block-$g$ is very similar. In particular both lead to higher power to detect small signal when compared with the $g$-prior. The tradeoff is a slightly higher (but acceptable) type I error.
@@
Figure 4: $F_{1}$ scores for model selection procedures based on various priors for our second simulation study.
@@
To further illustrate the false positive/false negative tradeoffs associated with the various methods, we present in Figure 4 boxplots across the various simulated dataset of the $F_{1}$ scores associated with the various procedures. As a reminder, the $F_{1}$ score is defined as the harmonic mean of proportion of true positives among “selected” covariates (the precision) and the proportion of “selected” covariates among true positive covariates (the recall). The $F_{1}$ score ranges between 0 and 1, with a higher value indicating better model selection performance. As would be expected from our previous discussion, ALasso is by far the worst performer, followed closely by Som et al. $(K=3)$. Som et al. $(K=2)$, GL-$g$ and DP block-$g$ are the best performers in all cases, with very minimal differences among them except perhaps for a small number of datasets in the most extreme scenario where $\eta=0.5$ and $p=750$. The $g$-prior is somewhere in between these two extremes.
Figure 5: Prediction MSE for $\eta=0$ and $\eta=0.5$.

### v2

@@ 4.6 Intrinsic consistency
$f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid a,b,\boldsymbol{\gamma})=\int\phi\left(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\boldsymbol{0},\sigma^{2}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{\Lambda}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right)\\ f\left(\tilde{\boldsymbol{g}}\mid\rho,\gamma,\tau^{2}=n,a,b\right)f(\rho\mid\boldsymbol{\gamma},\alpha)f(\alpha\mid\boldsymbol{\gamma})\mathop{}\!\mathrm{d}\tilde{g}_{1}\ldots\mathop{}\!\mathrm{d}\tilde{g}_{p_{\boldsymbol{\gamma}}}\mathop{}\!\mathrm{d}\rho\mathop{}\!\mathrm{d}\alpha.$
@@ 5 Computation
It is possible to construct MCMC algorithms for model selection under the DP mixtures of block $g$ prior that require very minimal tuning. To do so, we take advantage of the conditional conjugacy of the priors and, when possible, we integrate out the intercept $\beta_{0}$, the vector of regression coefficients $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ and/or the variance $\sigma^{2}$ when deriving conditional posteriors. Additionally, we represent the shrinkage coefficients $g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}}$ in terms of their unique values $\tilde{\boldsymbol{g}}_{\boldsymbol{\gamma}}=(\tilde{g}_{1},\ldots,\tilde{g}_{K_{\boldsymbol{\gamma}}})$ and the group indicators $\boldsymbol{\xi}_{\boldsymbol{\gamma}}=(\xi_{1},\ldots,\xi_{p_{\boldsymbol{\gamma}}})$ (recall Section 3). The resulting posterior takes the form
@@
where $f(\boldsymbol{y}\mid\boldsymbol{\gamma},\tilde{g}_{1},\ldots,\tilde{g}_{K_{\boldsymbol{\gamma}}},\xi_{1},\ldots,\xi_{p_{\boldsymbol{\gamma}}})$ corresponds to (3.2) with $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$, and $p(\boldsymbol{\gamma})$ is an appropriate prior on the space of models, e.g., a Beta-Binomial prior $f(\boldsymbol{\gamma})=\frac{\Gamma(c+d)}{\Gamma(c)\Gamma(d)}\frac{\Gamma(c+p_{\boldsymbol{\gamma}})\Gamma(d+p-p_{\boldsymbol{\gamma}})}{\Gamma(c+d+p)}$.
[…]
@@ 6.1 Conditional Lindley paradox
Our first simulation study replicates the setting used to construct Figure 1 and shows empirical evidence supporting the theoretical results discussed in Sections 4.3 and 4.4. We consider a total 150 simulations, each of which involves a sequence of datasets generated under model $\mathcal{M}_{a}:y_{i}=\beta_{0}+\beta_{1}x_{i,1}+\beta_{2}x_{i,2}+\epsilon_{i}$ for $i=1,\ldots,100$. All elements of a given sequence of datasets share the same values of $\beta_{0}$, $\beta_{1}$, $\boldsymbol{x}_{1}=(x_{1,1},\ldots,x_{100,1})^{\prime}$, $\boldsymbol{x}_{2}=(x_{1,2},\ldots,x_{100,2})^{\prime}$ and $\boldsymbol{\epsilon}=(\epsilon_{1},\ldots,\epsilon_{100})^{\prime}$. In particular, we set $\beta_{0}=0.5$, $\beta_{1}=1$ and generate $\epsilon$ from a standard multivariate Gaussian distribution and each of the pairs $(x_{i,1},x_{i,2})^{\prime}$ from a zero-mean bivariate Gaussian distribution with unit marginal standard deviations and correlation $\eta$. The datasets within each sequence are then constructed by considering a grid of values for $\beta_{2}$ in the interval $[0,240]$. We are interested in the behavior of $B_{a,0}$, the Bayes factor under the DP mixture of $g$ priors that uses the hyper-$g/n$ as the baseline measure comparing the true model $\mathcal{M}_{a}$ against the simpler model $\mathcal{M}_{0}:y_{i}=\beta_{0}+\beta_{1}x_{i,1}+\epsilon_{i}$. We also investigate the behavior of $\mathsf{Pr}(\xi_{1}\neq\xi_{2}\mid\boldsymbol{y})$, the posterior probability that the model assigns different shrinkage parameters to each variable in the model (recall Section 4.3).
Figure 2 shows the results of this simulation study for $\eta=0$ and $\eta=0.5$. Compared to Figure 1, the curves are somewhat noisy. This is an artifact of the Monte Carlo noise introduced by our MCMC algorithm (the Bayes factor depicted in 1 is available in “closed form”). With that caveat in mind, we note that the curves for $\log\left(B_{a,0}(\boldsymbol{y})\right)$ decrease as $\beta_{2}$ increases but, unlike Figure 1, both seem to stabilize towards an asymptote. This agrees with the behavior predicted by Theorem 4.4. Similarly, and as predicted by Theorem 4.3, we can see that $\mathsf{Pr}(\xi_{1}\neq\xi_{2}\mid\boldsymbol{y})$ seems to converge to 1 as $\beta_{2}$ grows.
@@ 6.2 Model selection, estimation and prediction performance
We conducted a second simulation study to compare the model selection, estimation, and prediction performance of procedures based on DP mixtures of block $g$ priors with that of competing procedures. We considered three versions of DP mixtures of block $g$ priors: one using a hyper-$g$ base measure (labeled “DP block-$g$ ($\tau^{2}=1$)”), a “unit information” version using a hyper-$g/n$ base measure (labeled “DP block-$g$ ($\tau^{2}=n$)”), and one using a scaled hyper-$g$ base measure where the scale is in turn given a half-Cauchy hyperprior (labeled “DP block-$g$ ($\tau^{2}\sim\mathsf{HC}$) ”). This second study, the setup of which is inspired by Denti et al. (2023), assumes $n=500$ for all datasets and, as before, the vectors of covariates associated with the each observation are generated from a zero-mean multivariate normal distribution with unit marginal variances and correlation $\eta$ across all pairs of covariates. We consider six scenarios that arise from combining three different values for the total number of covariates ($p=250$, $p=500$ and $p=750$) and two different levels of multicolinearity ($\eta=0$ and $\eta=0.9$). For all three values of $p$, 100100 of the coefficients are randomly sampled from a normal distribution with mean 0 and standard deviation 10 (we consider these “large” coefficients), $100$ are randomly sampled from a standard normal distribution (the “small” coefficients), and the remainder are set to 0 (the null coefficients). For each of the six scenarios, we generate 100 datasets.
@@
In terms of competing approaches, we consider the following: (a) a standard hyper-$g/n$ mixture of $g$ priors (which we label “g-prior” in the sequel); (b) a version of the block $g$ prior of Som (2014) with a hyper-$g/n$ hyperprior and known blocking structures where the covariates are allocated to $K=2$ groups: one made of all the covariates associated with large coefficients plus half, randomly chosen variables associated with null coefficients, and another one made of the rest (labeled “Som et al. ($\tau^{2}=n$, K=2K=2)”); (c) a version of the block $g$ prior of Som (2014) with $K=3$ fixed groups of covariates: one made of all the covariates associated with large coefficients, one made of the covariates associated with small coefficients, and one made of the covariates associated with null coefficients (labeled “Som et al. ($\tau^{2}=n$, K=3K=3)”); (d) three versions of “global-local” $g$-priors with distinct but identically distributed shrinkage parameters for each coefficient, one using a hyper-$g$ prior (labeled “GL-$g$ ($\tau^{2}=1$)”), one using a hyper-$g/n$ prior (labeled “GL-$g$ ($\tau^{2}=n$)”), and one using a scaled hyper-$g$ prior where the scale is given a half-Cauchy hyperprior (labeled “GL-$g$ ($\tau^{2}\sim\mathsf{HC}$) ”); (e) the adaptive Lasso (ALasso) of (Huang et al., 2008); (f) the Horseshoe prior (Carvalho et al., 2010, labeled “Horseshoe” in the sequel) and (g) the Horseshoe-Pit prior (Denti et al., 2023, labeled “HSM”). Computation under standard mixtures of $g$ priors relies on version 1.7.1 of the R package BAS, while computation under block $g$ priors with known blocking structures relies on a slight variation of our own code for the Diriclet mixtures of $g$ priors. Computation for the adaptive Lasso relies on version 4.1.6 of the R package glmnet. Computation under the Horseshoe prior relies on version 1.2 of the R package bayesreg, while computation under the Horseshoe-Pit prior relies on code from the author of that manuscript which, at the time of this writing, is available at https://github.com/Fradenti/HorseshoeMix. For all Bayesian procedures that require a prior on model space, we assign $\boldsymbol{\gamma}$ a Beta-Binomial prior with $c=d=1$ (e.g., see Scott and Berger, 2010 and Porwal and Raftery, 2022 and Section 5 above). Furthermore, in order to avoid improper priors when $p\geq n$, the prior on models is constrained so that models for which $\sum_{k=1}^{p}\gamma_{i}>n-2$ receive zero probability.
Power (large coeffs) Power (small coeffs) Type I error (null coeffs) Power (large coeffs) Power (small coeffs) Type I error (null coeffs)
@@
$\eta=0$ $\eta=0.9$
g-prior ($\tau^{2}=$n) 0.986 0.856 0.005 0.949 0.506 0.004
Som et al. ($\tau^{2}=$n, K=2) 0.986 0.913 0.037 0.962 0.850 0.219
Som et al. ($\tau^{2}=$n, K=3) 0.992 1.000 1.000 0.980 1.000 1.000
GL-$g$ ($\tau^{2}=$1) 1.000 1.000 1.000 1.000 1.000 1.000
GL-$g$ ($\tau^{2}=$n) 0.990 0.908 0.054 0.969 0.694 0.046
GL-$g$ ($\tau^{2}\simHC$) 0.990 0.900 0.039 0.973 0.745 0.100
DP block-g ($\tau^{2}=$1) 0.991 0.921 0.160 0.984 0.843 0.394
DP block-g ($\tau^{2}=$n) 0.989 0.904 0.045 0.974 0.760 0.126
DP block-g ($\tau^{2}\simHC$) 0.990 0.906 0.049 0.977 0.788 0.211
ALasso 0.956 0.569 0.011 0.873 0.352 0.157
Horseshoe 0.989 0.889 0.027 0.965 0.656 0.022
HSM 0.986 0.860 0.005 0.962 0.625 0.011
@@
$\eta=0$ $\eta=0.9$
g-prior ($\tau^{2}=$n) 0.979 0.814 0.000 0.913 0.342 0.001
Som et al. ($\tau^{2}=$n, K=2) 0.982 0.873 0.005 0.944 0.636 0.028
Som et al. ($\tau^{2}=$n, K=3) 0.986 0.923 1.000 0.963 0.843 1.000
GL-$g$ ($\tau^{2}=$1) 1.000 1.000 0.993 0.999 0.992 0.973
GL-$g$ ($\tau^{2}=$n) 0.984 0.866 0.006 0.955 0.580 0.010
GL-$g$ ($\tau^{2}\simHC$) 0.984 0.865 0.005 0.957 0.597 0.014
DP block-g ($\tau^{2}=$1) 0.985 0.873 0.014 0.961 0.636 0.040
DP block-g ($\tau^{2}=$n) 0.984 0.867 0.006 0.957 0.612 0.016
DP block-g ($\tau^{2}\simHC$) 0.984 0.867 0.007 0.958 0.622 0.023
ALasso 0.745 0.117 0.056 0.269 0.008 0.006
Horseshoe 0.982 0.842 0.012 0.947 0.540 0.007
HSM 0.981 0.831 0.001 0.943 0.506 0.001
@@
$\eta=0$ $\eta=0.9$
g-prior ($\tau^{2}=$n) 0.959 0.607 0.000 0.887 0.212 0.000
Som et al. ($\tau^{2}=$n, K=2) 0.982 0.863 0.003 0.941 0.576 0.015
Som et al. ($\tau^{2}=$n, K=3) 0.985 0.884 0.488 0.957 0.703 0.563
GL-$g$ ($\tau^{2}=$1) 0.992 0.922 0.468 0.976 0.792 0.464
GL-$g$ ($\tau^{2}=$n) 0.985 0.849 0.003 0.946 0.506 0.005
GL-$g$ ($\tau^{2}\simHC$) 0.985 0.851 0.003 0.948 0.526 0.008
DP block-g ($\tau^{2}=$1) 0.985 0.857 0.005 0.952 0.547 0.014
DP block-g ($\tau^{2}=$n) 0.985 0.852 0.003 0.949 0.536 0.010
DP block-g ($\tau^{2}\simHC$) 0.985 0.853 0.004 0.951 0.548 0.013
ALasso 0.722 0.106 0.054 0.242 0.007 0.004
Horseshoe 0.980 0.800 0.005 0.940 0.456 0.004
HSM 0.982 0.818 0.000 0.938 0.449 0.001
@@
Table 1: Estimates of power for small (generated from a $\mathcal{N}(0,1)$ distribution) and large (generated from a $\mathcal{N}(0,10)$ distribution) coefficients, and of type I error for null coefficients ($\beta=0$) in our second simulation study. For the purpose of this table, coefficients are considered “significant” is their posterior inclusion probability is greater than 0.5.
@@
In the introduction, we motivated DP mixtures of block $g$ priors by arguing that it should lead to higher power for detecting smaller coefficients. We also claimed that procedures that pre-select a blocking of the coefficients can be very sensitive to this choice. Evidence of these claims is presented in Table 1, which shows the estimated power associated with identifying “large” (those generated from a zero-mean normal with standard deviation 10) and “small” (those generated from a standard normal distribution) coefficients, as well as the type I error (for the null coefficients) under the various procedures. For the purpose of this table, coefficients are considered “significant” if their posterior inclusion probability is greater than 0.5 (in the case of procedures based on variants of the $g$-prior), if they are included in the optimal model after applying generalized cross-validation to identify the optimal penalty parameter (for ALasso), or if their 95% posterior credible intervals do not cover zero (in the case of Horseshoe and HSM). First, we note that ALasso shows by far the lowest power to detect small coefficients. Even for large coefficients, the performance of ALasso degrades substantially in “large pp” scenarios, especially when variables are highly correlated. The same is true of other penalized likelihood procedures we tried (results not shown). Next, we note that the standard $g$ prior tends to have lower power than the other Bayesian procedures we consider when it comes to detecting small coefficients. Indeed, somewhat surprisingly, HSM and Horseshoe seem to have higher overall power than standard $g$-priors in this scenario with similar type-I errors, even though they were not originally designed for model selection. GL-$g$ procedures with $\tau^{2}=1$ also seem to perform poorly, yielding high power for detecting small signals, but also very high type I error rates. Similarly to GL-$g$ with $\tau^{2}=1$, the procedure of Som (2014) with $K=3$ tends to have the highest power, but it also comes with extremely high type I error rates. This is because, by parceling out the null coefficients into a separate cluster, the block $g$ prior ends up overfitting by learning a very small shrinkage factor for the coefficients in this block. The results for Som (2014) with $K=2$ indicate that this issue can be addressed by having the null coefficients assigned to blocks that contain some significant coefficients, avoiding overfitting. This solution is, however, clearly impractical in real applications, as we do not know which coefficients are likely to be significant in the first place. The best performing procedures correspond to GL-$g$ ($\tau^{2}=n$), GL-$g$ ($\tau^{2}\sim\mathsf{HC}$), and the three procedures based on Dirichlet process mixtures of block $g$-priors. These are all priors that can be understood as being either “unit information”, or allowing for enough flexibility to learn the expected value of the $g_{j}$s from the data. For $\eta=0$, all of these 5 procedures seem to yield very similar performance in terms of power and type I errors. However, for $\eta=0.9$, procedures based on DP block-$g$ priors consistently show higher power for detecting small coefficients, at the price of a very small increase (in absolute terms) in the type I error.
@@
Figure 3: $F_{1}$ scores for model selection procedures for our second simulation study. These $F_{1}$ scores excludes the large coefficients in the calculation of precision and recall.
@@
To investigate the false positive/false negative tradeoffs associated with the various methods, we present in Figure 3 boxplots across the various simulated datasets of $F_{1}$ for the various procedures. Recall that the $F_{1}$ score is defined as the harmonic mean of proportion of true positives among “selected” covariates (the precision) and the proportion of “selected” covariates among true positive covariates (the recall). The $F_{1}$ score ranges between 0 and 1, with a higher value indicating better model selection performance. As would be expected from our discussion of power and false positive rates, ALasso is by far the worst performer, followed by Som et al. $(K=3)$ and procedures based on the GL-$g$ ($\tau^{2}=1$) prior. Procedures based on the standard $g$-prior perform somewhat better than these three, but they are still suboptimal, particularly when the covariates are highly correlated. The performance of the remaining procedures is very similar, but some subtle patterns are still visible. Som et al. $(K=2)$ seems to perform slightly better than the rest but, as discussed before, it relies on the unrealistic assumption that we are able to distinguish between large and small coefficients up front. Similarly, GL-$g$ ($\tau^{2}=n$), Horseshoe and HSM seem to slightly underperform, particularly when the correlation is high.
(a) $\eta=0$
(b) $\eta=0.9$
Figure 4: Normalized prediction MSE for $\eta=0$ and $\eta=0.9$. Normalization is with respect to the prediction MSE under the standard $g$-prior.